# Recommendations to OSCE/ODIHR (on how to give better recommendations for Internet voting)

### v1

@@ 3 Vote secrecy and coercion resistance
Another potential issue with vote secrecy is presented in the 2019 report making use of the fact that the Estonian Internet voting protocol implements individual verifiabiliy by making the encryption randomness available to the audit device via a QR-code [4]:
@@ 4 End-to-end verifiability
The intent of this recommendation is good, but it is not clear what is the correct level of technical detail such conditions should be written into the law. Invalidation of (part of) the voting results is a very prominent action, and the state may want to retain the discretion to decide flexibly based on the concrete circumstances that are hard to foresee in the time of writing the legislation. Thus, in 2017 an ammendment in the following wording entered force [^20] https://www.riigiteataja.ee/en/eli/ee/510032014001/consolide/current:
@@
Thus, the flexibility to consider the circumstances is there. On the other hand this of course means that it is impossible to give a strict mathematical proof that depends on the exact conditions for annulling the votes. [^21] Extending the argument we can see that this problem is inherent. The target of verification is detecting anomalities. In case of elections, the anomalities can hardly be fixed, so the only option is to cancel the voting event and try again. A precise mathematical definition of E2E verifiability hence has to rely on the (mathematical) criteria for annulling the result. In the real life, however, the annulling decision is taken by the lawyers who prefer not to be pre-committed to strict criteria, but rather have the power of discretion. It is ultimately the question of who gets to decide upon annulling the elections – mathematicians or lawyers. The current societal order seems to favour the lawyers in this role. Thus the mathematicians have to come up with E2E verifiabilty definitions that allow for the power of discretion for the lawyers, or to acknowledge that E2E verifiability can not be achieved in practice at all.

### v2

@@ 3 Vote secrecy and coercion resistance
Another potential issue with vote secrecy is presented in the 2019 report making use of the fact that the Estonian Internet voting protocol implements individual verifiability by making the encryption randomness available to the audit device via a QR-code [4]:
@@ 4 End-to-end verifiability
The intent of this recommendation is good, but it is not clear what is the correct level of technical detail such conditions should be written into the law. Invalidation of (part of) the voting results is a very prominent action, and the state may want to retain the discretion to decide flexibly based on the concrete circumstances that are hard to foresee in the time of writing the legislation. Thus, in 2017 an amendment in the following wording entered force [^20] https://www.riigiteataja.ee/en/eli/ee/510032014001/consolide/current:
@@
Thus, the flexibility to consider the circumstances is there. On the other hand this of course means that it is impossible to give a strict mathematical proof that depends on the exact conditions for annulling the votes. [^21] Extending the argument we can see that this problem is inherent. The target of verification is detecting anomalies. In case of elections, the anomalies can hardly be fixed, so the only option is to cancel the voting event and try again. A precise mathematical definition of E2E verifiability hence has to rely on the (mathematical) criteria for annulling the result. In the real life, however, the annulling decision is taken by the lawyers who prefer not to be pre-committed to strict criteria, but rather have the power of discretion. It is ultimately the question of who gets to decide upon annulling the elections – mathematicians or lawyers. The current societal order seems to favour the lawyers in this role. Thus the mathematicians have to come up with E2E verifiability definitions that allow for the power of discretion for the lawyers, or to acknowledge that E2E verifiability can not be achieved in practice at all.